# Connect Later: Improving Fine-tuning for Robustness with Targeted Augmentations

### v1

@@ 6 Discussion and Related Work
Data augmentations such as cropping or masking have been vital to semi- and self-supervised learning objectives. Masking or noising the data and training a model to reconstruct the original inputs have been shown to produce useful pretrained representations across multiple modalities [24, 35, 27, 36, 25, 26, 6]. In contrastive learning, models are trained to distinguish augmented “views” of the same input from views of a different input [25, 6, 26]. Our results demonstrating inconsistent OOD performance across datasets brings up the important future question of how to choose the best pretraining augmentation and algorithm for learning transferable representations.
@@
Data augmentation has been used to improve model robustness and avoid catastrophic failures due to spurious, label-independent changes (e.g. translation or rotation in vision) [22, 37, 38]. The augmentation strategies used in prior work are generic perturbations that aim to increase the diversity of inputs [e.g., 39, 40, 41, 42, 43, 44], though a number of studies have shown that the type of data augmentations matters for performance [25, 45]. Augmentations have also been leveraged in the self-training paradigm, which improves generalization to unseen data by training on the pseudo-labeled full dataset [34, 46, 47]. We show that a self-training baseline with pseudo-labels from an ERM+targeted augs model does not outperform Connect Later, indicating that pretraining is an important component of the framework. Connect Later exposes targeted augmentations as a design interface for improving robustness with knowledge of the distribution shift, while still leveraging pretrained representations.
@@
In problems with domain shift, Gao et al. [17] show that targeted augmentations outperform generic augmentations on unseen data. They identify spurious domain-dependent, label-independent features in the source dataset and construct targeted augmentations by randomizing these features. Gao et al. [17] consider the domain generalization setting, in which no data from the target dataset is available. We consider targeted augmentations in the domain adaptation setting, in which we can model the target distribution of these spurious features with the unlabeled target data. In general, designing targeted augmentations specific to each distribution shift may be difficult and require expert guidance. As part of the Connect Later framework, we provide a general methodology for the design of such augmentations. Certain aspects, such as the selection of feature space $z$ and transformation distribution $T$ could be learned from the unlabeled data itself, which we leave for future work. We show that targeted augmentations better leverage pretrained representations for complementary gains in OOD performance.
@@ 7 Conclusion
We show that pretraining with generic augmentations is not a panacea for all distribution shifts and tasks, and sometimes does not outperform supervised learning on labeled source data. Pure supervised learning, however, does not use the unlabeled data or produce reusable representations. The Connect Later framework allows for better leverage of pretrained representations for OOD performance by applying targeted augmentations at fine-tuning time. Future work could focus on learning targeted augmentations from the source and target data distributions as well as further understanding of how the choice of pretraining augmentations affects downstream ID/OOD performance.
@@ A.1 AstroClassification, Redshifts Datasets
The AstroClassification and Redshifts datasets were adapted from the 2019 Photometric LSST Astronomical Time-Series Classification Challenge [20]  [^3] https://zenodo.org/record/2539456. This diverse dataset contains 14 types of astronomical time-varying objects, simulated using the expected instrument characteristics and survey strategy of the upcoming Legacy Survey of Space and Time [LSST 48] conducted at the Vera C. Rubin Observatory. It includes two overall categories of time-series objects: transients, short-lived events such as supernovae, and variable sources, those with fluctuating brightness such as pulsating stars. Specifically, the dataset includes the following transients: type Ia supernovae (SNIa), SNIax, SNIa-91bg, SNIbc, SNII, superluminous supernovae (SLSN), tidal disruption events (TDE), and single lens microlensing events ($\mu$Lens-Single); and the following variable objects: active galactic nuclei (AGN), Mira variables, eclipsing binary systems (EB), and RR Lyrae (RRL).
@@
In these tasks, we are specifically interested in predicting the object type (e.g. type II supernova) and the cosmological redshift of objects in the unlabeled dataset. Cosmological redshift is a proxy for distance in the universe, and an important piece of metadata for understanding an object’s physical processes as well as other applications, such as estimating the expansion rate of the universe with type Ia supernovae.
@@
All data were simulated with the SuperNova ANAlysis [SNANA, 49] software library. Further details about the astrophysical models and LSST instrument characteristics used in the simulation can be found in Kessler et al. [50].
@@ B.1 Generic Augmentations for Pretraining
For the AstroClassification and Redshifts datasets, we randomly mask a subset of the input sequence using the masked language modeling paradigm introduced by [24]. Given an unlabeled input sequence $x$, a training input $x^{\prime}$ can be generated by randomly masking elements of $x$ while the associated label $y$ consists of the original, unmasked values. The model is trained to use contextual information (unmasked elements) to successfully reconstruct most of the sequence. From our ablation experiments, we find that a masking percentage of 60% produces the best downstream results. We follow an existing implementation for astronomical time-series [51] and set 80% of the masked elements to 0, replace 10% with a random element from the sequence, and keep the remaining 10% unchanged.
@@
For iWildCam-WILDS, we use a ResNet-50 model pretrained on ImageNet with SwAV, a contrastive learning algorithm [6]. For Camelyon17-WILDS, we use a DenseNet121 pretrained with SwAV on the unlabeled Camelyon17-WILDS dataset from Sagawa et al. [9]. SwAV uses random cropping augmentations of different resolutions.
@@ B.2 Targeted Augmentations for Fine-Tuning
The OOD test set of the AstroClassification and Redshifts datasets have many more high redshift objects than the source dataset, leading us to adopt an augmentation scheme to alleviate this shift. Figure 4 shows the redshift distributions of the source, augmented, and target datasets. Redshifting places each object at a new redshift and recomputes its light curve sampling, fluxes, and flux uncertainties accordingly. This augmentation algorithm was adapted from Boone [21].
@@
  • We model the observational capabilities of the telescope to ensure that our augmented input ${\bm{X}}^{\prime},{\bm{X}}^{\prime}_{\text{err}}$ does not fall below the threshold of detection. We “accept" an augmented input ${\bm{X}}^{\prime},{\bm{X}}^{\prime}_{\text{err}}$ if the signal-to-noise ratio (SNR) of at least two observations is over 5, i.e. $\text{SNR}({\bm{X}}^{\prime}_{i,j},{\bm{X}}^{\prime}_{\text{err},i,j})\geq 5$ for at least 2 of $i\in\{1,...,T\},j\in\{1,...,W\}$. We define $\text{SNR}(x,x_{\text{err}})=\frac{|x|}{x_{\text{err}}}$.
@@
This augmentation strategy randomizes the backgrounds of wildlife images to reduce the model’s dependence on these spurious features for species classification. Specifically, a segmentation mask is applied to each image to separate the animal from the background, and the animal is “copy-pasted" into a new background from a camera that has observed that animal species. This was the best performing augmentation strategy from Gao et al. [17].
@@
This augmentation, originally from Tellez et al. [29], alters the pixel values of the slide images to emulate different staining procedures used by different hospitals. The augmentation uses a pre-specified Optical Density (OD) matrix to project images from RGB space to a three-channel hematoxylin, eosin, and DAB space before applying a random linear combination. This was the best performing augmentation strategy from Gao et al. [17].
@@ Appendix C Experimental Details
We use an encoder-only Informer model [52] with 8 encoder layers of 12 attention heads each. The model hidden dimension was chosen to be 768 and the layer MLPs have hidden dimension 256. Due to the 2-dimensional position data (each element of the time-series has an associated time and photometric band/wavelength) and irregular sampling of our dataset, we train a positional encoding based on learnable Fourier features following Li et al. [53]. We also select a random window of length 300 from each example (and zero-pad examples with fewer than 300 observations) to produce inputs of uniform shape. We perform pretraining with a batch size of 256 and learning rate 1e-4 (selected from 1e-3 $\sim$ 1e-6) for 75,000 steps. We finetune the pretrained model with linear probing for 20,000 steps (for pretrained models only) and learning rate 1e-4, then fine-tuning for 10,000 steps at learning rate of 4e-5. We increase the learning rate for models without pretraining to 1e-4 for FT. The Redshifts task uses LP learning rate of 5e-4 and FT learning rate of 1e-4. We decrease the learning rate per step with a linear scheduler.
@@
For pretraining, we use ResNet-50 pretrained on ImageNet with SwAV [6]. During fine-tuning, we train all models for 15 epochs with early stopping on OOD validation performance, following Gao et al. [17]. For pretrained models, we also do 10 epochs of linear probing before fine-tuning [LP-FT, 28] for 15 epochs, where the linear probe is trained with Adam and the linear probe weights used to initialize the fine-tuning stage is chosen with OOD validation performance. To reduce the noise in OOD results, for all methods we select the epoch in the last 5 epochs with the best OOD validation performance and report OOD test results with that version of the model. Following Gao et al. [17], we allow for 10 hyperparameter tuning runs, where we sample the following hyperparameters independently from the following distributions: the linear probe learning rate ($10^{\text{Uniform}[-3,-2]}$), fine-tuning learning rate ($10^{\text{Uniform}[-5,-2]}$), and probability of applying the augmentation ($\text{Uniform}[0.5,0.9]$) and pick the hyperparameter configuration with the best OOD validation performance. For ERM and ERM+targeted augmentations, we use the tuned hyperparameters from Gao et al. [17]. To decrease the confidence interval due to an outlier seed, the reported performance of Connect Later is averaged over 15 seeds. All other results are averaged over 5 seeds.
@@
For pretraining, we use DenseNet121 pretrained on the unlabeled Camelyon17-WILDS dataset presented in Sagawa et al. [9] with SwAV [6]. During fine-tuning, we train all models for 15 epochs with early stopping on OOD validation performance, following Gao et al. [17]. For pretrained models, we also do 10 epochs of linear probing before fine-tuning [LP-FT, 28] for 15 epochs, where the linear probe is trained with Adam and the linear probe weights used to initialize the fine-tuning stage is chosen with OOD validation performance. To reduce the noise in OOD results, for all methods we select the epoch with the best OOD validation performance and report OOD test results with that version of the model. Following Gao et al. [17], we allow for 10 hyperparameter tuning runs, where we sample the following hyperparameters independently from the following distributions: the linear probe learning rate ($10^{\text{Uniform}[-3,-2]}$), fine-tuning learning rate ($10^{\text{Uniform}[-5,-2]}$), probability of applying the augmentation ($\text{Uniform}[0.5,0.9]$), and augmentation strength ($\text{Uniform}[0.05,0.1]$), and pick the hyperparameter configuration with the best OOD validation performance. All results are averaged over 20 seeds.
@@ Appendix D Empirical Estimates of Connectivity
We empirically estimate connectivity measures for all of the datasets we tested on following the procedure outlined in Appendix D of Shen et al. [7]. Specifically, we train binary classifiers from scratch to predict the class-domain pair of a given input example. We randomly select 15 class-domain pairs for iWildCam-WILDS and AstroClassification, while for Camelyon17-WILDS we use all class-domain pairs since Camelyon17-WILDS is a binary classification task. We label these class-domain examples following Appendix D of Shen et al. [7] and create a dataset with 80/10/10 train/validation/test split. We train using the same hyperparameters described in Appendix C for 3,000 steps with early stopping on the validation accuracy. Our results are presented in Table 7.
@@ Appendix E Simple construction where Connect Later improves over pretraining or targeted augmentations alone
We give a simple construction for constrastive pretraining based on the construction in Proposition 3 (Appendix A.2) of Shen et al. [7], where Connect Later improves over pretraining (standard fine-tuning) or targeted augmentations alone.
@@
Following HaoChen et al. [54] and Shen et al. [7], we analyze contrastive learning from an augmentation graph perspective, where inputs $x$ are connected via augmentations with edge weights $S_{+}(x,{x^{\prime}})$, which represent the probability of $x,{x^{\prime}}$ being a positive pair (augmentations of the same input $x$). For theoretical analysis, we analyze the spectral contrastive learning objective:
@@
Figure 5: Example distribution of data and augmentations for contrastive learning where Connect Later improves OOD performance over contrastive pretraining+standard fine-tuning and ERM+targeted augmentations. The augmentation graph is similar to Shen et al. [7] except the edge weights connecting 1,2 and 3,4 are swapped. The shapes represent classes, while the labeled data is shaded in green. The generic augmentation probabilities are marked as edge weights, where we assume that $\alpha>\gamma+\beta$. Here, targeted augmentations which first swap inputs 1 and 2 before applying a generic augmentation help to align the source and target. However, some target inputs are not reachable via augmentations from source inputs. Standard fine-tuning can generalize throughout the target domain, but only in conjunction with targeted augmentations that align the source and target. The orange dotted lines on the far ends connect to each other (the graph wraps around).
@@
Instead of analyzing fine-tuning, we follow Shen et al. [7] and analyze linear probing on top of the pretrained representations from the encoder. We train a linear model with parameters $B\in\mathbb{R}^{r\times k}$, where $r$ is the number of classes. We minimize the objective:
@@
Notice that the weight between 1,3 is ${\gamma^{\prime}}$ and the weight between 1,4 is ${\alpha^{\prime}}$, and the weights are similarly swapped for 2,4, and 2,5. We assume that ${\rho^{\prime}},{\alpha^{\prime}},{\beta^{\prime}}$, and ${\gamma^{\prime}}$ are in $(0,1)$ and are distinct. We also assume that the augmentation probabilities satisfy ${\rho^{\prime}}>\max\{{\alpha^{\prime}},{\beta^{\prime}}\}$ and $\min\{{\alpha^{\prime}},{\beta^{\prime}}\}>{\gamma^{\prime}}$. Following Shen et al. [7], we can convert these to positive pair probabilities ${\rho},{\alpha},{\beta},{\gamma}$ with similar properties by renormalizing.
Given the above setting, the following is a simplified form of Proposition 3 from Shen et al. [7], if we instead use the following augmentation distribution, which swaps the edge weight magnitudes that involve nodes 1 and 2:
@@
Proposition 1 (Shen et al. [7]).
[…]
@@
In general, we proceed by defining the following targeted augmentation, which allows us to reduce to the setting of Proposition 1:
@@
By Proposition 1, standard fine-tuning after contrastive pretraining has zero target (OOD) error when the pretraining augmentations do not have swapped edges. By symmetry, standard fine-tuning (contrastive pretraining + linear probing) on our augmentation graph with pretraining augmentations $\mathcal{A}_{\text{pre}}$ outputs the opposite label for all target inputs, resulting in an OOD error of 1. This is because the source and target domains are misaligned in our augmentation graph.
@@
Connect Later applies targeted augmentations $\mathcal{A}_{\text{ft}}$ during the linear probing step (on top of contrastive pretrained representations). This choice of targeted augmentations reduces to the setting of Proposition 1 where the labeled source domain consists of the inputs 3,4 instead. By the symmetry of the graph and applying Proposition 1, Connect Later achieves 0 OOD error.

### v2

@@ 6 Discussion and Related Work
Data augmentations (e.g., cropping or masking) are vital to semi- and self-supervised learning objectives. Reconstructing a masked or noised input has been shown to produce useful pretrained representations across multiple modalities (Devlin et al., 2019; Lewis et al., 2020; He et al., 2022; Raffel et al., 2019; Chen et al., 2020; He et al., 2020; Caron et al., 2020). In contrastive learning, models are trained to distinguish augmented “views” of the same input from views of a different input (Chen et al., 2020; Caron et al., 2020; He et al., 2020). Our results demonstrate that though pretraining with generic augmentations alone produces inconsistent OOD performance across datasets, fine-tuning with targeted augmentations is able to better leverage these pretrained representations.
@@
Data augmentation has been used to improve model robustness to label-independent changes (e.g. translation or rotation in vision) (Hendrycks et al., 2019; Rebuffi et al., 2021; Ng et al., 2020). Existing augmentation strategies rely on generic perturbations that aim to increase the diversity of inputs (e.g., Simard et al., 2003; Krizhevsky et al., 2012; Cubuk et al., 2019, 2020; DeVries & Taylor, 2017; Zhang et al., 2017), though prior work has shown that the type of data augmentations matters for performance (Chen et al., 2020; Xie et al., 2020a). Augmentations have also been leveraged in the self-training paradigm, which improves generalization to unseen data by training on the pseudo-labeled full dataset (Xie et al., 2020b; Sohn et al., 2020; Yang et al., 2021). We show that a self-training baseline with pseudo-labels from an ERM+targeted augs model does not outperform Connect Later, indicating that pretraining is important for robustness gains. Connect Later exposes targeted augmentations as a design interface for improving robustness with knowledge of the distribution shift, while still leveraging pretrained representations.
@@
In domain shift problems, Gao et al. (2023) show that targeted augmentations designed with knowledge of the distribution shift outperform generic, or even target-aware (e.g. CutMix, Yun et al. (2019)), augmentations on unseen data. Gao et al. (2023) consider the domain generalization setting, in which the target dataset is unknown. We consider targeted augmentations in the domain adaptation setting, where we can model the target distribution with the unlabeled target data. In this setting, targeted augmentations provide the opportunity to naturally incorporate domain knowledge about the dataset and distribution shift. In this work, we provide a general methodology for the design of such augmentations and show that targeted augmentations better leverage pretrained representations for complementary gains in OOD performance. Certain aspects of the design process, such as the selection of feature space $z$ and transformation distribution $T$ could be learned from the unlabeled data itself, which we leave for future work.
@@ 7 Conclusion
We show that pretraining with generic augmentations is not a panacea for all distribution shifts and tasks, and does not deliver consistent gains over supervised learning on labeled source data. Pure supervised learning, however, does not use the unlabeled data or produce reusable representations. Connect Later allows for better leverage of pretrained representations for OOD performance by applying targeted augmentations at fine-tuning time.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Acknowledgments
We thank the anonymous reviewers for their comments, leading to substantial improvements of the paper. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award HEP-dessn.
@@ A.1 AstroClassification, Redshifts Datasets
The AstroClassification and Redshifts datasets were adapted from the 2019 Photometric LSST Astronomical Time-Series Classification Challenge (The PLAsTiCC team et al., 2018)  [^2] https://zenodo.org/record/2539456. This diverse dataset contains 14 types of astronomical time-varying objects, simulated using the expected instrument characteristics and survey strategy of the upcoming Legacy Survey of Space and Time (LSST Ivezić et al., 2019) conducted at the Vera C. Rubin Observatory. It includes two overall categories of time-series objects: transients, short-lived events such as supernovae, and variable sources, those with fluctuating brightness such as pulsating stars. Specifically, the dataset includes the following transients: type Ia supernovae (SNIa), SNIax, SNIa-91bg, SNIbc, SNII, superluminous supernovae (SLSN), tidal disruption events (TDE), and single lens microlensing events ($\mu$Lens-Single); and the following variable objects: active galactic nuclei (AGN), Mira variables, eclipsing binary systems (EB), and RR Lyrae (RRL).
@@
In these tasks, we are specifically interested in predicting the object type (e.g. type II supernova) and the cosmological redshift of objects in the unlabeled dataset. Cosmological redshift is a proxy for distance in the universe, and an important piece of metadata for understanding an object’s physical processes as well as other applications, such as estimating the expansion rate of the universe with type Ia supernovae. The redshift prediction task has been studied for individual object types, such as quasars (Nakoneczny et al., 2021) and type Ia supernovae (Qu & Sako, 2023), but we consider a more realistic set of multiple object types.
@@
All data were simulated with the SuperNova ANAlysis (SNANA, Kessler et al., 2009) software library. Further details about the astrophysical models and LSST instrument characteristics used in the simulation can be found in Kessler et al. (2019).
@@ B.1 Generic Augmentations for Pretraining
For the AstroClassification and Redshifts datasets, we randomly mask a subset of the input sequence using the masked language modeling paradigm introduced by (Devlin et al., 2019). Given an unlabeled input sequence $x$, a training input $x^{\prime}$ can be generated by randomly masking elements of $x$ while the associated label $y$ consists of the original, unmasked values. The model is trained to use contextual information (unmasked elements) to successfully reconstruct most of the sequence. From our ablation experiments, we find that a masking percentage of 60% produces the best downstream results. We follow an existing implementation for astronomical time-series (Donoso-Oliva et al., 2023) and set 80% of the masked elements to 0, replace 10% with a random element from the sequence, and keep the remaining 10% unchanged.
@@
For iWildCam-WILDS, we use a ResNet-50 model pretrained on ImageNet with SwAV, a contrastive learning algorithm (Caron et al., 2020). For Camelyon17-WILDS, we use a DenseNet121 pretrained with SwAV on the unlabeled Camelyon17-WILDS dataset from Sagawa et al. (2022). SwAV uses random cropping augmentations of different resolutions.
@@ B.2 Targeted Augmentations for Fine-Tuning
The OOD test set of the AstroClassification and Redshifts datasets have many more high redshift objects than the source dataset, leading us to adopt an augmentation scheme to alleviate this shift. Figure 4 shows the redshift distributions of the source, augmented, and target datasets. Redshifting places each object at a new redshift and recomputes its light curve sampling, fluxes, and flux uncertainties accordingly. This augmentation algorithm was adapted from Boone (2019).
@@
  • We model the observational capabilities of the telescope to ensure that our augmented input ${\bm{X}}^{\prime},{\bm{X}}^{\prime}_{\text{err}}$ does not fall below the threshold of detection. We “accept” an augmented input ${\bm{X}}^{\prime},{\bm{X}}^{\prime}_{\text{err}}$ if the signal-to-noise ratio (SNR) of at least two observations is over 5, i.e. $\text{SNR}({\bm{X}}^{\prime}_{i,j},{\bm{X}}^{\prime}_{\text{err},i,j})\geq 5$ for at least 2 of $i\in\{1,...,T\},j\in\{1,...,W\}$. We define $\text{SNR}(x,x_{\text{err}})=\frac{|x|}{x_{\text{err}}}$.
@@
This augmentation strategy randomizes the backgrounds of wildlife images to reduce the model’s dependence on these spurious features for species classification. Specifically, a segmentation mask is applied to each image to separate the animal from the background, and the animal is “copy-pasted” into a new background from a camera that has observed that animal species. This was the best performing augmentation strategy from Gao et al. (2023).
@@
This augmentation, originally from Tellez et al. (2018), alters the pixel values of the slide images to emulate different staining procedures used by different hospitals. The augmentation uses a pre-specified Optical Density (OD) matrix to project images from RGB space to a three-channel hematoxylin, eosin, and DAB space before applying a random linear combination. This was the best performing augmentation strategy from Gao et al. (2023).
@@ Appendix C Experimental Details
We use an encoder-only Informer model (Zhou et al., 2021) with 8 encoder layers of 12 attention heads each. The model hidden dimension was chosen to be 768 and the layer MLPs have hidden dimension 256. Due to the 2-dimensional position data (each element of the time-series has an associated time and photometric band/wavelength) and irregular sampling of our dataset, we train a positional encoding based on learnable Fourier features following Li et al. (2021). We also select a random window of length 300 from each example (and zero-pad examples with fewer than 300 observations) to produce inputs of uniform shape. We perform pretraining with a batch size of 256 and learning rate 1e-4 (selected from 1e-3 $\sim$ 1e-6) for 75,000 steps. We finetune the pretrained model with linear probing for 20,000 steps (for pretrained models only) and learning rate 1e-4, then fine-tuning for 10,000 steps at learning rate of 4e-5. We increase the learning rate for models without pretraining to 1e-4 for FT. The Redshifts task uses LP learning rate of 5e-4 and FT learning rate of 1e-4. We decrease the learning rate per step with a linear scheduler.
@@
For pretraining, we use ResNet-50 pretrained on ImageNet with SwAV (Caron et al., 2020). During fine-tuning, we train all models for 15 epochs with early stopping on OOD validation performance, following Gao et al. (2023). For pretrained models, we also do 10 epochs of linear probing before fine-tuning (LP-FT, Kumar et al., 2022) for 15 epochs, where the linear probe is trained with Adam and the linear probe weights used to initialize the fine-tuning stage is chosen with OOD validation performance. To reduce the noise in OOD results, for all methods we select the epoch in the last 5 epochs with the best OOD validation performance and report OOD test results with that version of the model. Following Gao et al. (2023), we allow for 10 hyperparameter tuning runs, where we sample the following hyperparameters independently from the following distributions: the linear probe learning rate ($10^{\text{Uniform}[-3,-2]}$), fine-tuning learning rate ($10^{\text{Uniform}[-5,-2]}$), and probability of applying the augmentation ($\text{Uniform}[0.5,0.9]$) and pick the hyperparameter configuration with the best OOD validation performance. For ERM and ERM+targeted augmentations, we use the tuned hyperparameters from Gao et al. (2023). To decrease the confidence interval due to an outlier seed, the reported performance of Connect Later is averaged over 15 seeds. All other results are averaged over 5 seeds.
@@
For pretraining, we use DenseNet121 pretrained on the unlabeled Camelyon17-WILDS dataset presented in Sagawa et al. (2022) with SwAV (Caron et al., 2020). During fine-tuning, we train all models for 15 epochs with early stopping on OOD validation performance, following Gao et al. (2023). For pretrained models, we also do 10 epochs of linear probing before fine-tuning (LP-FT, Kumar et al., 2022) for 15 epochs, where the linear probe is trained with Adam and the linear probe weights used to initialize the fine-tuning stage is chosen with OOD validation performance. To reduce the noise in OOD results, for all methods we select the epoch with the best OOD validation performance and report OOD test results with that version of the model. Following Gao et al. (2023), we allow for 10 hyperparameter tuning runs, where we sample the following hyperparameters independently from the following distributions: the linear probe learning rate ($10^{\text{Uniform}[-3,-2]}$), fine-tuning learning rate ($10^{\text{Uniform}[-5,-2]}$), probability of applying the augmentation ($\text{Uniform}[0.5,0.9]$), and augmentation strength ($\text{Uniform}[0.05,0.1]$), and pick the hyperparameter configuration with the best OOD validation performance. All results are averaged over 20 seeds.
@@ Appendix D Empirical Estimates of Connectivity
We empirically estimate connectivity measures for all of the datasets we tested on following the procedure outlined in Appendix D of Shen et al. (2022). Specifically, we train binary classifiers from scratch to predict the class-domain pair of a given input example. We randomly select 15 class-domain pairs for iWildCam-WILDS and AstroClassification, while for Camelyon17-WILDS we use all class-domain pairs since Camelyon17-WILDS is a binary classification task. We label these class-domain examples following Appendix D of Shen et al. (2022) and create a dataset with 80/10/10 train/validation/test split. We train using the same hyperparameters described in Appendix C for 3,000 steps with early stopping on the validation accuracy. Our results are presented in Table 8.
@@ Appendix E Simple construction where Connect Later improves over pretraining or targeted augmentations alone
We give a simple construction for constrastive pretraining based on the construction in Proposition 3 (Appendix A.2) of Shen et al. (2022), where Connect Later improves over pretraining (standard fine-tuning) or targeted augmentations alone.
@@
Following HaoChen et al. (2021) and Shen et al. (2022), we analyze contrastive learning from an augmentation graph perspective, where inputs $x$ are connected via augmentations with edge weights $S_{+}(x,{x^{\prime}})$, which represent the probability of $x,{x^{\prime}}$ being a positive pair (augmentations of the same input $x$). For theoretical analysis, we analyze the spectral contrastive learning objective:
@@
Figure 5: Example distribution of data and augmentations for contrastive learning where Connect Later improves OOD performance over contrastive pretraining+standard fine-tuning and ERM+targeted augmentations. The augmentation graph is similar to Shen et al. (2022) except the edge weights connecting 1,2 and 3,4 are swapped. The shapes represent classes, while the labeled data is shaded in green. The generic augmentation probabilities are marked as edge weights, where we assume that $\alpha>\gamma+\beta$. Here, targeted augmentations which first swap inputs 1 and 2 before applying a generic augmentation help to align the source and target. However, some target inputs are not reachable via augmentations from source inputs. Standard fine-tuning can generalize throughout the target domain, but only in conjunction with targeted augmentations that align the source and target. The orange dotted lines on the far ends connect to each other (the graph wraps around).
@@
Instead of analyzing fine-tuning, we follow Shen et al. (2022) and analyze linear probing on top of the pretrained representations from the encoder. We train a linear model with parameters $B\in\mathbb{R}^{r\times k}$, where $r$ is the number of classes. We minimize the objective:
@@
Notice that the weight between 1,3 is ${\gamma^{\prime}}$ and the weight between 1,4 is ${\alpha^{\prime}}$, and the weights are similarly swapped for 2,4, and 2,5. We assume that ${\rho^{\prime}},{\alpha^{\prime}},{\beta^{\prime}}$, and ${\gamma^{\prime}}$ are in $(0,1)$ and are distinct. We also assume that the augmentation probabilities satisfy ${\rho^{\prime}}>\max\{{\alpha^{\prime}},{\beta^{\prime}}\}$ and $\min\{{\alpha^{\prime}},{\beta^{\prime}}\}>{\gamma^{\prime}}$. Following Shen et al. (2022), we can convert these to positive pair probabilities ${\rho},{\alpha},{\beta},{\gamma}$ with similar properties by renormalizing.
Given the above setting, the following is a simplified form of Proposition 3 from Shen et al. (2022), if we instead use the following augmentation distribution, which swaps the edge weight magnitudes that involve nodes 1 and 2:
@@
Proposition E.1 (Shen et al. (2022)).
With the above construction for the input space $\mathcal{X}$, unlabeled distribution $P_{U}$, and data augmentation $\mathcal{A}_{\text{prop}}$, for some feature dimension $k\in\mathbb{Z}^{+}$ a linear probe trained on contrastive pre-trained features achieves 0 target error: $\mathcal{L}_{0-1}(\widehat{f})=0$. However, for all $k\in\mathbb{Z}^{+}$, there exists a minimizer $\widehat{f}_{\text{erm}}$ of the ERM objective (with data augmentations according to $\mathcal{A}_{\text{prop}}$) that has non-zero error: $\mathcal{L}_{0-1}(\widehat{f}_{\text{erm}})=1/3$.
@@
In general, we proceed by defining the following targeted augmentation, which allows us to reduce to the setting of Proposition E.1:
@@
By Proposition E.1, standard fine-tuning after contrastive pretraining has zero target (OOD) error when the pretraining augmentations do not have swapped edges. By symmetry, standard fine-tuning (contrastive pretraining + linear probing) on our augmentation graph with pretraining augmentations $\mathcal{A}_{\text{pre}}$ outputs the opposite label for all target inputs, resulting in an OOD error of 1. This is because the source and target domains are misaligned in our augmentation graph.
@@
Connect Later applies targeted augmentations $\mathcal{A}_{\text{ft}}$ during the linear probing step (on top of contrastive pretrained representations). This choice of targeted augmentations reduces to the setting of Proposition E.1 where the labeled source domain consists of the inputs 3,4 instead. By the symmetry of the graph and applying Proposition E.1, Connect Later achieves 0 OOD error.